# Grounding Synthetic Data Evaluations of Language Models in Unsupervised Document Corpora

### v2

@@ Abstract
Language Models (LMs) continue to advance, improving response quality and coherence. Given Internet-scale training datasets, LMs have likely encountered much of what users may ask them to generate in some form during their training. A plethora of evaluation benchmarks have been constructed to assess model quality, response appropriateness, and reasoning capabilities. However, the human effort required for benchmark construction is rapidly being outpaced by the size and scope of the models under evaluation. Having humans build a benchmark for every possible domain of interest is impractical. Therefore, we propose a methodology for automating the construction of fact-based synthetic data model evaluations grounded in document populations. This work leverages the same LMs to evaluate domain-specific knowledge automatically, using only grounding documents (e.g., a textbook) as input. This synthetic data benchmarking approach corresponds well with human curated questions producing a Spearman ranking correlation of and a benchmark evaluation Pearson accuracy correlation of . This novel approach supports generating both multiple choice and open-ended synthetic data questions to gain diagnostic insight of LM capability. We apply this methodology to evaluate model performance on two recent arXiv preprints, discovering a surprisingly strong performance from Gemma-3 models on open-ended questions. Code is available at https://github.com/mmajurski/grounded-synth-lm-benchmark
@@ 1 Introduction
Recent advances in language models (LMs) have demonstrated significant capability improvements in diverse natural language processing tasks. This expansion in capability has primarily been enabled by training on Internet-scale corpora. As AI tooling improves professional workflows, rigorous evaluation is imperative to understand the domain-specific model capability.
Traditional multiple choice question (MCQ) benchmarks such as MMLU [1] offer valuable general insights into model performance, but are static, public, and limited in scope. The fact that these benchmarks are static and public—and therefore can appear in training data—is a significant weakness. Combined with the human effort required to construct new high quality tests [2], it is unlikely evaluation construction can keep pace with LM development. This scaling challenge is especially acute for domain-specific professional applications, where the general benchmarks provide limited insight. Unfortunately, creating expert-validated benchmarks for every domain is impractical. New automated methods must generate accurate, comprehensive, diagnostically relevant evaluations grounded in authoritative corpora with minimal human intervention beyond document selection.
@@
Do these synthetic benchmarks replicate existing human-curated evaluations?
@@ 2 Related works
Our approach is motivated by task asymmetry similar to NP-completeness: given grounding documents, an LM should reliably generate diagnostically relevant questions, while zero-shot answering the same questions is substantially harder. This information asymmetry motivates using LMs to generate benchmarks grounded in curated documents rather than relying only on model weight knowledge. Language models are routinely adopted as judges: extracting answer choices [3, 4], grading open ended responses [5, 6], and justifying correctness [7]. The most similar contemporary convergent work is YourBench [8] synthetic replication of MMLU from grounding documents and AutoBencher [9] reinforcement learning optimization to discover gaps in LM knowledge. Leveraging grounding documents greatly improves synthetic data utility, as the methodology no longer relies on internalized LM knowledge.
Modern long-context LMs can process entire documents. Benchmarks like SecQA [10] and PubMedQA [11] leverage domain-specific textbooks and web pages. Full-document grounding generally outperforms summary-based systems [12]. Continued LM improvement, prompting strategies [13], and topic breadth [14] has further improved synthetic data. These methods also enhance downstream applications like retrieval-augmented generation [15].
LMs can serve as acceptable feature extractors when guided by detailed prompts. Studies show strong correlation between LM-based and human judgments of text quality [16, 17]. This aligns with the LM-as-a-Judge paradigm used for open-ended benchmark evaluation [18, 19]. Beyond open ended answer correctness, LMs can extract other textual features. However, limitations in LM ground truthing remain, especially in tasks like information retrieval where relevance judgments are more complex [20, 21]. LM evaluation benchmarks generally use either multiple choice (MC) or open-ended (OE) formats. Open-ended questions require synthesis [22] but depend on a judge LM for scoring [18, 23], while MCQs are easier to grade but may inflate performance due to answer visibility [24, 25]. Frameworks like Inspect-AI [26] support both types with prompt templates and answer shuffling. Both formats are sensitive to rewording, leading to inconsistent outputs [27], and newer frameworks aim to improve open-ended scoring precision [28, 29].
@@ 3 Methods
Constructing synthetic data benchmarks from relevant domain-specific documents requires chunking into contexts, topic extraction per context, question and answer generation, and finally an evaluation framework to determine model capability on the newly minted benchmark. All LM interactions use default generation parameters, except we set (o4-mini did not support setting temperature).
@@
While modern LMs have sufficient context windows to support whole document ingestion, chunking encourages the generated questions to be spread throughout the grounding document. An input document is first converted into markdown for interoperability with LMs. We use Docling [30] for conversion from pdf to markdown, and pandoc [31] for conversion from html to markdown. Once in markdown, each document section is used as a context chunk for question generation. This chunking method is simple and relies on the document author crafting relevant sections. Enhancements to chunking and document data preparation will likely produce overall system improvements, but were not required to demonstrate baseline performance in this study. For instance, our approach only leverages the textual data; formatting and including richer sources of information from table, charts, and images may improve evaluation quality.
@@
Given (a) the document context chunk and (b) a single topic for a question, the generating LM is prompted to construct a high quality and difficult evaluation question to probe the knowledge and understanding of an expert in the domain. During question generation, the LM is required to write the question, correct answer, and a justification/explanation for why the correct answer is in fact right. The answer justifications are not used in the evaluation benchmark and simply exist to guard against generating incorrect answers. In summary, the single grounding document is broken into chunks of context, per-context relevant topics are extracted, then per-context/topic pair, and a single evaluation question is created. This results in a list of questions with correct answers grounded in (and derived from) the source document. This example drawn from Squadv2 [32] shows a question generated by Llama-3.3-70B-Instruct grounded in the context:
@@ 3.1 Benchmark evaluation
Given a list of questions and answers, an LM evaluation framework (Inspect-AI [26]) is used to determine performance for the models under evaluation. The benchmark consists of many questions with associated answers. The resulting benchmark evaluation accuracy is the average number of correct answers divided by the total number of questions. LMs are evaluated against both the synthetic evaluation benchmark and the original questions. The overall accuracy per evaluated model can be compared for both the synthetic and original questions across multiple datasets of grounding documents. The text of the various benchmark questions and answers is never directly compared, only the resulting LM performance. Figure 2 showcases the performance of synthetic multiple choice questions averaged across datasets and generating models, with a Spearman ranking correlation of , demonstrating slight model over-performance on the synthetic benchmark.
@@ 4 Evaluation methodology
Evaluation of the quality and utility of the newly minted synthetic data benchmark built from grounding documents is done by comparing the synthetic data pipeline on datasets where human annotated questions and correct answers exist for known context paragraphs. This provides a directly comparable result, both for the overall benchmark accuracy (between human and grounded synthetic data) and for the rank ordering of those models that the benchmark is evaluated on. We leverage six Natural Language Processing (NLP) datasets where each entry contains a context paragraph, human written question, and correct answer. For each dataset a variety of LMs are used to generate the synthetic benchmark. Per-question quality features and benchmark-level question diversity is extracted to compare against the human written questions. Finally, an automatic quality control pass is completed to grade both question and answer correctness/validity, in order to investigate and characterize which benchmark-generating models can write valid question and answer pairs.
@@ 4.1 Datasets
Evaluating the validity of generated synthetic questions requires source documents which have human annotated question/answer pairs to compare against. The following six NLP datasets were selected: (a) Squadv2 [32] (b) HotpotQA [33] (c) TrivaQA-web [34] (d) NaturalQuestionsShort [35] (e) PubMedQA [11], and (f) SecQA [10].
@@
| If Roman numerals were used, what would Super Bowl 50 have been called? | Super Bowl L | What would the 50th Super Bowl have been called under the traditional Roman-numeral naming? | (A) Super Bowl XL (B) Super Bowl XLVIII (C) Super Bowl XLIX (D) Super Bowl L |
These datasets collectively contain tens of thousands of context, question, answer tuples. A random subset of 100 tuples per dataset was sampled as the grounding context document. For certain datasets, pre-processing was required to clarify, curate, and reformat the source human written questions into self-contained questions. For some questions this required disambiguation so the question is answerable without the context. For the multiple choice evaluation, three believable incorrect answer options needed to be created. The PubMedQA [11] dataset required conversion from yes/no/maybe questions into multiple-choice or open-ended. Only the SecQA [10] dataset already contained all of the required information and was not pre-processed. All other datasets went through an LM based reformatting operation correct these deficiencies. Table 1 provides an example of the reformatting operation for a few questions drawn from a single context of the Squadv2 dataset.
Three different models (Llama-4-Maverick-17B-128E-Instruct, gemma-3-27b-it, o4-mini) were used for this reformatting operation. The remainder of this paper shows results based on the Llama-4-Maverick-17B-128E-Instruct reformatting (see Table 4 in Appendix E for results of all three reformatting models).
@@ 4.2 Question generation models
| gemma-3-4b-it [36] | |
| gemma-3-12b-it [36] | |
| gemma-3-27b-it [36] | |
| Meta | Llama-3.2-3B-Instruct [37] |
| Llama-3.1-8B-Instruct [38] | |
| Llama-3.3-70B-Instruct [38] | |
| Llama-4-Maverick-17B-128E-Instruct [39] | |
| OpenAI | o4-mini [40] |
| gpt-4.1 [40] | |
| gpt-4.1-mini [40] | |
| Microsoft | phi-4 [41] |
| Qwen | Qwen2.5-7B-Instruct [42] |
To explore the impact of model capability on successful generation of synthetic data benchmarks we tested a variety of language models during the synthetic benchmark generation phase (see Table 2). All open-weight models are available through HuggingFace [43] and the closed-weight models were accessed through paid APIs. Additional results are reported for an ensemble of the bolded models. During benchmark creation. the same model is used both for topic extraction and question generation.
@@ 4.4 Generated question quality characterization
Per-question quality features and benchmark dataset-level features are computed to determine the overall quality of the synthetic benchmark. Owing to the low level of human oversight and annotation time for creating the synthetic data benchmarking, LMs extract these somewhat imprecise features. This relies on the idea that with enough guidance LMs can be good alternatives to human evaluation [16, 17] and borrows from standard practice LLM-as-a-Judge grading of open-ended benchmark responses [18, 19]. Llama-4-Maverick-17B-128E-Instruct extracts the following information about each generated question:
@@ 5 Results
All experiments and development were conducted on a single 8×H100 on-premise machine. Development consumed approximately twice the compute of final evaluations. Final compute was split between synthetic dataset generation and benchmark evaluation via Inspect-AI. The most expensive model, gpt-4.1, was only used during final results, processing 3.6M input and 1.1M output tokens ( $16), with compute evenly divided between dataset creation and evaluation.
@@ 5.1 Answer correctness
Synthetic benchmark answers must be accurate, a task made easier via the grounding documents that supply the necessary information. To verify generated answer accuracy, an LM scored each question-answer pair and its justification (1-10 scale) using the full context. As shown in Figure 3, reasonably sized models scored above 9.5 in accuracy, with smaller models performing noticeably worse. While model size/capability generally correlates with desirable properties for synthetic data generation, notable exceptions (like gpt-4.1) are discussed later.
@@ 5.2 Question clarity and groundedness
How grounded in the source material (context chunk) a given question is indicates how much outside information both the generation and evaluation models must leverage to correctly interpret what is being asked and answer. Low grounding indicates the question/answer pair cannot be found within the context chunk, or that significant external information is required. High grounding means all necessary information is found within the context chunk. Therefore the generated question is testing for information aligned with and present in the context. Figure 4 highlights the question clarity and groundedness scores for the question generating models. While most LMs generate similar clarity values, the trends in groundedness score are more complex. Some capable models (like Llama-4-Maverick-17B-128E-Instruct) score well, other capable models (like o4-mini) score worse, with smaller models in between (gemma-3-4b-it). We observed two grounding failure modes: either the model hallucinates irrelevant content absent from the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). See Appendix F for an example from phi-4. Models of medium size perform well on groundedness, simply reformatting information presented in the context.

### v3

# Grounding Generative Evaluations of Language Models in Unsupervised Document Corpora

@@ Abstract
Language Models (LMs) continue to advance, improving response quality and coherence. Given Internet-scale training datasets, LMs have likely encountered much of what users may ask them to generate in some form during their training. A plethora of evaluation benchmarks have been constructed to assess model quality, response appropriateness, and reasoning capabilities. However, the human effort required for benchmark construction is rapidly being outpaced by the size and scope of the models under evaluation. Having humans build a benchmark for every possible domain of interest is impractical. Therefore, we propose a methodology for automating the construction of fact-based synthetic data model evaluations grounded in document populations. This work leverages the same LMs to evaluate domain-specific knowledge automatically, using only grounding documents (e.g., a textbook) as input. This generative benchmarking approach corresponds well with human curated questions producing an ensemble Spearman ranking correlation of and a benchmark evaluation Pearson accuracy correlation of (model specific ). This novel approach supports generating both multiple choice and open-ended synthetic data questions to gain diagnostic insight of LM capability. We apply this methodology to evaluate model performance on three recent documents (two post LM knowledge cutoff), discovering a surprisingly strong performance from Gemma-3 models on open-ended questions. Code is available at https://github.com/mmajurski/grounded-synth-lm-benchmark
@@ 1 Introduction
Recent advances in language models (LMs) have demonstrated significant capability improvements in diverse natural language processing tasks. This expansion in capability has primarily been enabled by training on Internet-scale corpora. While LMs generate fluent, plausible text, the outputs are not always grounded in the facts of a situation. As AI tooling improves professional workflows, rigorous evaluation is imperative to understand the domain-specific model capability.
Traditional multiple choice question (MCQ) benchmarks such as MMLU (Hendrycks et al., 2020) offer valuable general insights into model performance, but are static, public, and limited in scope. The fact that these benchmarks are static and public—and therefore can appear in training data—is a significant weakness. Combined with the human effort required to construct new high quality tests (Phan et al., 2025), it is unlikely evaluation construction can keep pace with LM development. This scaling challenge is especially acute for domain-specific professional applications, where the general benchmarks provide limited insight. Unfortunately, creating expert-validated benchmarks for every domain is impractical. New automated methods must generate accurate, comprehensive, diagnostically relevant evaluations grounded in authoritative corpora with minimal human intervention beyond document selection.
@@
Do these generative benchmarks replicate existing human-curated evaluations?
@@ 2 Related works
Our approach is motivated by task asymmetry where given grounding documents, an LM should reliably generate diagnostically relevant questions, while it is substantially harder to zero-shot answer those same questions. This information asymmetry motivates using LMs to generate benchmarks grounded in curated documents rather than relying only on model weight knowledge. Language models are routinely adopted as judges: extracting answer choices (Samuylova, 2024; Roucher, 2024), grading open ended responses (Ip, 2024; Oh et al., 2024), and justifying correctness (Cook et al., 2024). The most similar contemporary convergent work is YourBench (Shashidhar et al., 2025) synthetic replication of MMLU from grounding documents and AutoBencher (Li et al., 2025b) reinforcement learning optimization to discover gaps in LM knowledge. Leveraging grounding documents greatly improves synthetic data utility, as the methodology no longer relies on internalized LM knowledge.
Modern long-context LMs can process entire documents. Benchmarks like SecQA (Liu, 2023) and PubMedQA (Jin et al., 2019) leverage domain-specific textbooks and web pages. Full-document grounding generally outperforms summary-based systems (Bhat et al., 2023). Continued LM improvement, prompting strategies (Yadav et al., 2024), and topic breadth (Li et al., 2024) has further improved synthetic data. These methods also enhance downstream applications like retrieval-augmented generation (Balaguer et al., 2024).
LMs can serve as acceptable feature extractors when guided by detailed prompts. Studies show strong correlation between LM-based and human judgments of text quality (Chiang and Lee, 2023; Han et al., 2024). This aligns with the LM-as-a-Judge paradigm used for open-ended benchmark evaluation (Xu et al., 2023; Zheng et al., 2023). Beyond open ended answer correctness, LMs can extract other textual features. However, limitations in LM ground truthing remain, especially in tasks like information retrieval where relevance judgments are more complex (Soboroff, 2025; Szymanski et al., 2025). LM evaluation benchmarks generally use either multiple choice (MC) or open-ended (OE) formats. Open-ended questions require synthesis (Balepur et al., 2025) but depend on a judge LM for scoring (Xu et al., 2023; Myrzakhan et al., 2024), while MCQs are easier to grade but may inflate performance due to answer visibility (Robinson et al., 2022; Balepur et al., 2024). Frameworks like Inspect-AI (UK AI Security Institute, 2024) support both types with prompt templates and answer shuffling. Both formats are sensitive to rewording, leading to inconsistent outputs (Wang et al., 2024a), and newer frameworks aim to improve open-ended scoring precision (Bernard et al., 2024; Kamalloo et al., 2023).
@@ 3 Methods
Constructing synthetic data benchmarks from relevant domain-specific documents requires chunking into contexts, topic extraction per context, question and answer generation, and finally an evaluation framework to determine model capability on the newly minted benchmark (as shown in Figure˜2). All LM interactions use default generation parameters, except we set (o4-mini did not support setting temperature).
@@
While modern LMs have sufficient context windows to support whole document ingestion, chunking encourages the generated questions to be spread throughout the grounding document. An input document is first converted into markdown for interoperability with LMs. We use Docling (Auer et al., 2024) for conversion from pdf to markdown, and pandoc (MacFarlane, ) for conversion from html to markdown. Once in markdown, each document section is used as a context chunk for question generation. This chunking method is simple and relies on the document author crafting relevant sections. Enhancements to chunking and document data preparation will likely produce overall system improvements, but were not required to demonstrate baseline performance in this study. For instance, our approach only leverages the textual data; formatting and including richer sources of information from table, charts, and images may improve evaluation quality.
@@
Given (a) the document context chunk and (b) a single topic for a question, the generating LM is prompted to construct a high quality and difficult evaluation question to probe the knowledge and understanding of an expert in the domain. During question generation, the LM is required to write the question, correct answer, and a justification/explanation for why the correct answer is in fact right. The answer justifications are not used in the evaluation benchmark and simply exist to guard against generating incorrect answers. In summary, the single grounding document is broken into chunks of context, per-context relevant topics are extracted, then per-context/topic pair, and a single evaluation question is created. This results in a list of questions with correct answers grounded in (and derived from) the source document. This example drawn from Squadv2 (Rajpurkar et al., 2018) shows a question generated by Llama-3.3-70B-Instruct grounded in the context:
@@ 3.1 Benchmark evaluation
Given a list of questions and answers, an LM evaluation framework (Inspect-AI (UK AI Security Institute, 2024)) is used to determine performance for the models under evaluation. The benchmark consists of many questions with associated answers. Open ended questions are graded using the default Inspect-AI framework support which uses a narrowly scoped LM-Judge to compare a candidate model output to the ground truth answer to determine correctness. The resulting benchmark evaluation accuracy is the average number of correct answers divided by the total number of questions. LMs are evaluated against both the synthetic evaluation benchmark and the original extractive QA human written questions lightly rewritten to be answerable without the context. The overall accuracy per evaluated model can be compared for both the synthetic and original questions across multiple datasets of grounding documents. The text of the various benchmark questions and answers is never directly compared, only the resulting LM performance. Figure˜3 showcases the performance of synthetic multiple choice questions averaged across datasets and generating models, with a Spearman ranking correlation of , demonstrating slight model over-performance on the synthetic benchmark.
@@ 4 Evaluation methodology
Evaluating the quality and utility of the newly minted synthetic data benchmark built upon grounding documents leveraged Natural Language Processing (NLP) datasets where each entry contains a context paragraph, human written question, and correct answer. This provides a directly comparable result, both for the overall benchmark accuracy (between human and grounded synthetic data) and for the rank ordering of those models that the benchmark is evaluated on. For each dataset a variety of LMs are used to generate the synthetic benchmark. Per-question quality features and benchmark-level question diversity is extracted to compare against the human written questions. Finally, an automatic quality control pass is completed to grade both question and answer correctness/validity, in order to investigate and characterize which benchmark-generating models can write valid question and answer pairs.
@@ 4.1 Datasets
Evaluating the validity of generated synthetic questions requires source documents which have human annotated question/answer pairs to compare against. The following NLP datasets were evaluated: (a) Squadv2 (Rajpurkar et al., 2018) (b) HotpotQA (Yang et al., 2018) (c) TrivaQA-web (Joshi et al., 2017) (d) NaturalQuestionsShort (Kwiatkowski et al., 2019) (e) PubMedQA (Jin et al., 2019) (f) BoolQ (Clark et al., 2019) (g) FermiQA (Kalyan et al., 2021) (h) MS-MARCO-QA (Bajaj et al., 2016) (i) MusiqueQA (Trivedi et al., 2022) (j) NarrativeQA (Kočiskỳ et al., 2018) (k) 2WikiMultiHopQA (Ho et al., 2020), and (l) SecQA (Liu, 2023).
@@
| If Roman numerals were used, what would Super Bowl 50 have been called? | Super Bowl L | What would the 50th Super Bowl have been called under the traditional Roman‐numeral naming? | (A) Super Bowl XL (B) Super Bowl XLVIII (C) Super Bowl XLIX (D) Super Bowl L |
These datasets collectively contain tens of thousands of context, question, answer tuples. A random subset of 100 samples per dataset was selected as the grounding context document. For certain datasets, pre-processing was required to clarify, curate, and reformat the source human written questions into self-contained questions. For some questions this required disambiguation so the question is answerable without the context. For the multiple choice evaluation, three believable incorrect answer options needed to be created. The PubMedQA (Jin et al., 2019) and BoolQ (Clark et al., 2019) datasets required conversion from yes/no questions into multiple-choice or open-ended. Only the SecQA (Liu, 2023) dataset already contained all of the required information and was not pre-processed. All other datasets went through an LM based reformatting operation correct these deficiencies. Table˜1 provides an example of the reformatting operation for a few questions drawn from a single context of the Squadv2 dataset. This disambiguation rewrite was validated using both an LM-Judge and embedding cosine similarity to ensure the questions were not unduly altered. Throughout the rest of this work, original questions refers to these lightly edited and disambiguated version of the human authored questions.
| gemma-3-4b-it (Team et al., 2025) | |
| gemma-3-12b-it (Team et al., 2025) | |
| gemma-3-27b-it (Team et al., 2025) | |
| Meta | Llama-3.2-3B-Instruct (36) |
| Llama-3.1-8B-Instruct (Grattafiori et al., 2024) | |
| Llama-3.3-70B-Instruct (Grattafiori et al., 2024) | |
| Llama-4-Maverick-17B-128E-Instruct (58) | |
| OpenAI | o4-mini (42) |
| gpt-4.1 (42) | |
| gpt-4.1-mini (42) | |
| Microsoft | phi-4 (Abdin et al., 2024) |
| Qwen | Qwen2.5-7B-Instruct (Yang et al., 2024) |
| Qwen3-235B-A22B-Instruct-2507 (Team, 2025) |
Three different models (Llama-4-Maverick, gemma-3-27b-it, o4-mini) were evaluated for this reformatting operation. The remainder of this paper shows the Llama-4-Maverick reformatting (see Table˜4 in Appendix˜F for the minimally different results of other reformatting models).
@@ 4.2 Question generation models
To explore the impact of model capability on successful generation of synthetic data benchmarks we tested a variety of language models during the synthetic benchmark generation phase (see Table˜2). All open-weight models are available through HuggingFace (39) and the closed-weight models were accessed through paid APIs. Additional results are reported for an ensemble of the bolded models where all generated questions from model ensemble were mixed together as an evaluation benchmark. During benchmark creation, the same model is used both for topic extraction and question generation.
@@ 4.4 Generated question quality characterization
Per-question quality features and benchmark dataset-level features are computed to determine the overall quality of the synthetic benchmark. Owing to the low level of human oversight and annotation time for creating the synthetic data benchmarking, LMs extract these somewhat imprecise features. This relies on the idea that with enough guidance LMs can be good alternatives to human evaluation (Chiang and Lee, 2023; Han et al., 2024) and borrows from standard practice LLM-as-a-Judge grading of open-ended benchmark responses (Xu et al., 2023; Zheng et al., 2023). Llama-4-Maverick extracts the following information about each generated question:
@@ 5 Results
All experiments and development were conducted on a single 8×H100 on-premise machine. Development consumed approximately twice the compute of final evaluations. Final compute was split between synthetic dataset generation and benchmark evaluation via Inspect-AI. The most expensive model, gpt-4.1 ingests about 3.6M input and 1.1M output tokens for an average evaluation run ( $16), with compute evenly divided between generative dataset creation and evaluation.
@@ 5.1 Answer correctness
Synthetic benchmark answers must be accurate, a task made easier via the grounding documents that supply the necessary information. To verify generated answer accuracy, an LM scored each question-answer pair and its justification (1-10 scale) using the full context as described in Section˜4.4. As shown in Figure˜4, reasonably sized models scored above 9.5 in accuracy, with smaller models performing noticeably worse. While model size/capability generally correlates with desirable properties for synthetic data generation, notable exceptions (like gpt-4.1-mini) are discussed in Figure 14 of Appendix F.
@@ 5.2 Question clarity and groundedness
How grounded in the source material (context chunk) a given question is indicates how much outside information both the generation and evaluation models must leverage to correctly interpret what is being asked and answer. Low grounding indicates the question/answer pair cannot be found within the context chunk, or that significant external information is required. High grounding means all necessary information is found within the context chunk. Therefore the generated question is testing for information aligned with and present in the context. Figure˜5 highlights the question clarity and groundedness scores for the question generating models. While most LMs generate similar clarity values, the trends in groundedness score are more complex. Some capable models (like Llama-4-Maverick) score well, other capable models (like o4-mini) score worse, with smaller models in between (gemma-3-4b-it). We observed two grounding failure modes: either the model hallucinates irrelevant content absent from the context chunk (phi-4 commonly did this), or the generating model uses knowledge in its weights to expand upon the information presented in the context (o4-mini). See Appendix˜G for an example from phi-4. Models of medium size perform well on groundedness, simply reformatting information presented in the context.